OpenStudy (supersquidsisters):

help? when i graphed this the graph turned out really wild, and i dont know how else to solve it Where are the asymptotes of f(x) = tan(4x − π) from x = 0 to x = pi over 2 ? (1 point) x = pi over 4 , x = 3 pi over 4 x = 0, x = pi over 4 x = pi over 2 , x = 3 pi over 2 x = 3 pi over 8 , x = 5 pi over 8

satellite73 (satellite73):

asympotes are where the input is equal to \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) solve this by solving \[4x-\pi=-\frac{\pi}{2}\] and \[4x-\pi=\frac{\pi}{2}\]

satellite73 (satellite73):

steps to solve both are the same a) add \(\pi\) to both sides then b) divide both sides by \(4\)

satellite73 (satellite73):

here is a nice picture if you need one http://www.wolframalpha.com/input/?i=tan(4x-pi)+domain+0..pi%2F2

OpenStudy (supersquidsisters):

@satellite73 ok so for the first one i got 1.78539.... and .39269..... are these right??

satellite73 (satellite73):

i would leave the answers with pi's in them, since those are your choices

OpenStudy (supersquidsisters):

oh ok so how would it be written as a fraction?

satellite73 (satellite73):

lets take it slow

satellite73 (satellite73):

\[4x-\pi=-\frac{\pi}{2}\]solve for \(x\) step one add \(\pi\) ti both sides

satellite73 (satellite73):

\[4x=\pi-\frac{\pi}{2}\] so i guess the question is, what is \(\pi-\frac{\pi}{2}\) do you know it ? (no is a fine answer, just asking)

OpenStudy (supersquidsisters):

i know as a decimal it would be 1.57079.....

satellite73 (satellite73):

i meant as a fraction

satellite73 (satellite73):

here is an easier question, what is \[1-\frac{1}{2}\]

OpenStudy (supersquidsisters):

1/2, right?

satellite73 (satellite73):

yes, right

satellite73 (satellite73):

so by the same arithmetic \[\pi-\frac{\pi}{2}=\frac{\pi}{2}\]

satellite73 (satellite73):

hmm now that i look at your answer choices, i am not sure i like any of them, so i wonder what is going on allow me to solve the first one for you, \[4x-\pi=-\frac{\pi}{2}\] first add \(\pi\) to get \[4x=-\frac{\pi}{2}+\pi=\frac{\pi}{2}\] then divide by \(4\) to get \[x=\frac{\pi}{8}\]

satellite73 (satellite73):

but that is not one of your answer choices, so lets try \[4x-\pi=\frac{\pi}{2}\] add \(\pi\) get \[4x=\frac{\pi}{2}+\pi=\frac{3\pi}{2}\]

satellite73 (satellite73):

divide by \(4\) get \(x=\frac{3\pi}{8}\)

satellite73 (satellite73):

which at least has the benefit of being one of your choices i wonder what happened to \(\frac{\pi}{8}\)?

OpenStudy (supersquidsisters):

would that be the 5pi over 8???

satellite73 (satellite73):

i guess it you solve \[4x-\pi=\frac{3\pi}{2}\] you get it , yes

satellite73 (satellite73):

question has an error in it , don't fret about it, not worth your effort pick the last one and move on

OpenStudy (supersquidsisters):

oh, ok!! thank you for all your help!!! do you think you can also help me with another one im having trouble with?

satellite73 (satellite73):

sure no problem always happy to help a squid

OpenStudy (supersquidsisters):

ok!! ill post it as a separate question so i can give you another medal for it

satellite73 (satellite73):

not necessary, i have plenty

satellite73 (satellite73):

but you can if you like